OpenStudy (anonymous):

If you have a piece of paper that is 0.1 mm thick, how many times will you have to fold it in half in order for it to become tall enough reach the moon?

OpenStudy (anonymous):

@Legends @Nnesha

OpenStudy (misty1212):

hi!!

OpenStudy (misty1212):

\[0.1\times 2^n=distance\]

OpenStudy (misty1212):

you need the distance to the moon

OpenStudy (legends):

A giant leap to hood

OpenStudy (legends):

to man hood

OpenStudy (anonymous):

aww nnesha is doing another problem whatever. lol

OpenStudy (misty1212):

it varies, but you can use 15 billion inches

OpenStudy (anonymous):

Nope @legends thats not right :D

OpenStudy (legends):

Am I close

OpenStudy (anonymous):

You guys give up yet?

OpenStudy (misty1212):

so \[0.1\times 2^n=15,000,000,000\] solve for \(n\)

OpenStudy (misty1212):

or \[2^n=150,000,000,000\] answer is \[\frac{\ln(150,000,000,000)}{\ln(2)}\]

OpenStudy (anonymous):

The fold doubles every time you fold it, after 10 folds it's over 10 cm thick. SO after 42 folds, its around 439804 km thick, tall enough to reach the moon. :D

OpenStudy (anonymous):

there are alot of tricky math problems here. http://time.com/3108579/tricky-math-problems-quiz/